2x+7y=4 and y=2x+5...Are they Parallel,perpendicular,or neither?
Found 2 solutions by Gogonati, ewatrrr:
Answer by Gogonati(855)

These lines are neither parallel or perpendicular.

Answer by ewatrrr(24785)

 
Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
2x+7y=4 Or y = (-2/7)x + 4/7 m = (-2/7)
y=2x+5 m = 2
Are they Parallel,perpendicular,or neither? Neither
parallel have identical slopes
perpendicular have slopes that are negative reciproclas of one another:
ex: m -2/7 and m = 7/2